OpenStudy (anonymous):

Point A(–4, 2) is reflected over the line x = 3 to create the point A'. What are the coordinates of A'?

OpenStudy (anonymous):

what kind of line is x = 3?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

x = 3 is a straight line (vertical), crossing the x-axis at x = 3.

OpenStudy (anonymous):

If it is reflected, then the distance between the point A and the line x = 3 and A' and the line x = 3 has to be equal!

OpenStudy (anonymous):

So the distance in x from A x to the line x = 3 is 3-(-4) = 7. The distance in y does not change, cause the line x = 3 is only vertical! Therefore, the y coordinates of A and A' are the same, but the x coordinates differ. We found the distance in x to be 7 between x = 3 and A, then it also has to be for x = 3 and A', which leads us to the conclusion that the x coordinate of A' must be x = 3 + 7 = 10. Hence, A'(10, 2)

OpenStudy (anonymous):

*So the distance in x from A to...